Bonaime, Inc., has 7.5 million shares of common stock outstanding. The current share price is $62.50, and the book value per share is $5.50. The company also has two bond issues outstanding. The first bond issue has a face value of $71.5 million, a coupon rate of 7 percent, and sells for 90.5 percent of par. The second issue has a face value of $36.5 million, a coupon rate of 8 percent, and sells for 89.5 percent of par. The first issue matures in 20 years, the second in 12 years.

a. What are the company’s capital structure weights on a book value basis? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)
  

Book value weight of equity
Book value weight of debt


b. What are the company’s capital structure weights on a market value basis? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)
  

Market value weight of equity
Market value weight of debt


c. Which are more relevant, the book or market value weights?
  
(Click to select) Market value weights Book value weights
Answer #1

a

BV of equity=Price of equity*number of shares outstanding
BV of equity=5.5*7500000
=41250000
BV of Bond1=Par value*bonds outstanding*%age of par
BV of Bond1=1000*71500*1
=71500000
BV of Bond2=Par value*bonds outstanding*%age of par
BV of Bond2=1000*36500*1
=36500000
BV of firm = BV of Equity + BV of Bond1+ BV of Bond 2
=41250000+71500000+36500000
=149250000
Weight of equity = BV of Equity/BV of firm
Weight of equity = 41250000/149250000
W(E)=0.2764
Weight of debt = BV of Bond/BV of firm
Weight of debt = 108000000/149250000
W(D)=0.7236

b

MV of equity=Price of equity*number of shares outstanding
MV of equity=62.5*7500000
=468750000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*71500*0.905
=64707500
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*36500*0.895
=32667500
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=468750000+64707500+32667500
=566125000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 468750000/566125000
W(E)=0.828
Weight of debt = MV of Bond/MV of firm
Weight of debt = 97375000/566125000
W(D)=0.172

c

market value weights are more relevant as the cost of capital for debt and equity are based due to the market value of the sources of capital